OpenStudy (anonymous):

the square root of 97 is between what 2 integers

OpenStudy (rowanmartinfan):

ILL HELP U

OpenStudy (rowanmartinfan):

BTWN 96 & 98

OpenStudy (anonymous):

Note that \(\sqrt{81}<\sqrt{97}<\sqrt{100}\). How do the lower and upper bounds simplify? :-)

OpenStudy (rowanmartinfan):

HE ISNT LOOKING @ OUR CHATING